# Machine Learning and the Physical Sciences (ML4PS)

### v1

@@ 1 Introduction
Oscillator Ising Machines (OIMs) are natural candidates for EP because their dynamics follow energy gradient descent [2]. Unlike discrete Ising solvers [8], OIMs have continuous phase dynamics that support the precise nudging required by EP, with couplings and biases acting as trainable parameters. Originally developed to minimize Ising energies, they can now be repurposed as neuromorphic machine learning processors.
The synergy between EP and OIMs offers the prospect of fast, energy-efficient neuromorphic computing. GPU training consumes substantial power [4], while OIMs achieve both speed and efficiency by exploiting GHz-frequency physical dynamics for rapid energy descent [3]. This acceleration could make energy-based learning practical, overcoming the long relaxation and sampling times that limit such models on conventional processors [1].
Previous EP implementations on oscillators faced initialization or synchronization issues [14, 11]. We present the first framework that maps EP directly onto existing OIM hardware without modification, turning devices originally designed for combinatorial optimization into neural network processors. This illustrates a broader principle: when physical dynamics coincide with ML optimization, they can be directly harnessed to build faster and more efficient learning systems.
@@ 2 Oscillator Ising Machines
An Oscillator Ising Machine (OIM) consists of a network of $n$ coupled self-sustaining nonlinear oscillators with non-varying amplitudes and equal frequencies. Over a sufficiently coarse-grained timescale [15], each oscillator can be parameterized by a phase $\phi_{i}\in[0,2\pi]$ whose dynamics follow $\frac{d\phi_{i}}{dt^{\prime}}=-\sum_{j=1,\ j\neq i}^{n}J_{ij}\sin(\phi_{i}-\phi_{j})-h_{i}\sin(\phi_{i})-S_{i}\sin(2\phi_{i})$.
@@
These dynamics are equivalent to gradient descent $\frac{d\phi_{i}}{dt^{\prime}}=-\frac{\partial V}{\partial\phi_{i}}$, for an energy function [^1] Symmetric $J_{ij}$ includes $\frac{1}{2}$ to avoid double counting. [2]: $V=-\frac{1}{2}\sum_{i,j\ i\neq j}^{n}J_{ij}\cos(\phi_{i}-\phi_{j})-\sum_{i}^{n}h_{i}\cos(\phi_{i})-\sum_{i}^{n}\frac{S_{i}}{2}\cos(2\phi_{i}).$
@@ 3 Equilibrium Propagation
Equilibrium Propagation (EP) [13, 7] is a local alternative to backpropagation for networks whose dynamics are characterized by the convergence to a stationary point, avoiding a separate backward pass by nudging these steady states toward desired outputs.
@@
The updates to the trainable parameters $\{\theta\}$ are then given by [7] $\Delta\theta=\eta\hat{\nabla}^{\rm EP}(\beta)$ for learning rate $\eta$ and where:
@@
This update effectively steers the free stationary point toward configurations with lower loss by decreasing energy at $s^{\beta}_{*}$ and increasing it at $s^{-\beta}_{*}$. In the limit of infinitesimal $\beta$, this update equals gradient descent on the loss [7]:
@@
Moreover, during the nudged phase dynamics (before convergence), the instantaneous EP updates match Backpropagation Through Time (BPTT) in the $\beta\to 0$ limit [16]:
@@ 4 Implementation on Oscillator Ising Machines
We implement a dense neural network on an OIM with $n_{x}$ input neurons (fixed during dynamics), $n_{h}$ hidden neurons, and $n_{y}$ output neurons. Each non-input neuron corresponds to an oscillator, with the dynamical state $\phi=(\phi^{(h)},\phi^{(y)})$, and trainable parameters ${\theta}=({w_{ij}^{(x,h)}},{w_{ij}^{(h,y)}},{b_{i}^{(h)}},{b_{i}^{(y)}})$.
@@
Training follows standard EP. We use He weight initialization[6] with zero biases. For each input $x$: (i) Configure the OIM parameters as specified above; (ii) Initialize all oscillators to the reference state [^2] This corresponds to setting all neuron activations $\cos(\frac{\pi}{2})=0$. We found that any random initialization of phases also works well, provided the same random reference state is used consistently across all dynamics. $\phi_{0}=\{\frac{\pi}{2}\}$; (iii) Run the free phase ($\beta=0$) dynamics until convergence to $\phi_{*}$; (iv) Run the positive and negative nudged phase dynamics (initialized at $\phi_{*}$) until convergence to $\phi_{*}^{+\beta}$ and $\phi_{*}^{-\beta}$.
@@ 5 Results
We validate our OIM–EP implementation on MNIST [9], Fashion-MNIST [17], and the MNIST/100 subset (1,000/100 train/test used in prior Ising machine benchmarks [8]). Training used PyTorch with Euler integration of the OIM dynamics. Hyperparameters (integration steps and size, $\beta$, learning rates, batch size) were set conventionally and tuned for stability.
@@
OIM networks achieve 90.6±1.7% on MNIST/100 (92.0±0.3% with noise), and 97.2±0.1%/88.0±0.1% on full MNIST/FMNIST (Fig. 1). These results surpass D-Wave annealers [8] and match BPTT on the same OIMs (96.8±0.1% on MNIST), indicating accuracy is architecture-limited rather than training-limited. On FMNIST, OIMs outperform p-bit Ising machines [10] and are competitive with other neuromorphic MLPs [18].
@@
Overall, OIMs maintain accuracy under 10-bit parameters, 4-bit phase detection [5], and moderate phase noise, confirming viability for CMOS implementations [12].
@@ 6 Discussion
Physical OIMs promise dramatic speedups. With GHz-frequency oscillators reaching microsecond timescales for parameter updates and phase measurements [15], full MNIST training (50 epochs, 60k examples, 3 EP phases) could complete in seconds to minutes, compared to $\sim$40 hours for EP and $\sim$60 hours for BPTT in our simulations—an acceleration of several orders of magnitude. This speed and efficiency could make energy-based learning practical for tasks previously limited by training time.

### v2

# Learning at the Speed of Physics: Equilibrium Propagation on Oscillator Ising Machines

@@ 1 Introduction
Oscillator Ising Machines (OIMs) are natural candidates for EP because their dynamics follow energy gradient descent [basharNoteAnalyzingStability2023]. Unlike discrete Ising solvers [laydevantTrainingIsingMachine2024], OIMs have continuous phase dynamics that support the precise nudging required by EP, with couplings and biases acting as trainable parameters. Originally developed to minimize Ising energies, they can now be repurposed as neuromorphic machine learning processors.
The synergy between EP and OIMs offers the prospect of fast, energy-efficient neuromorphic computing. GPU training consumes substantial power [garcia-martinEstimationEnergyConsumption2019], while OIMs achieve both speed and energy efficiency by exploiting GHz-frequency physical dynamics for rapid energy descent [csabaCoupledOscillatorsComputing2020b]. This acceleration could make energy-based learning practical, overcoming the long relaxation and sampling times that limit such models on conventional processors [assranSelfSupervisedLearningImages2023].
Previous EP implementations on oscillators faced initialization or synchronization issues [wangTrainingCoupledPhase2024, rageauTrainingSynchronizingOscillator2025]. We present the first framework that maps EP directly onto existing OIM hardware without modification, turning devices originally designed for combinatorial optimization into neural network processors; a more detailed write-up of this study is available in [gowerOIMEPFull]. This illustrates a broader principle: when physical dynamics coincide with ML optimization, they can be directly harnessed to build faster and more efficient learning systems.
@@ 2 Oscillator Ising Machines
An Oscillator Ising Machine (OIM) consists of a network of $n$ coupled self-sustaining nonlinear oscillators with non-varying amplitudes and equal frequencies. Over a sufficiently coarse-grained timescale [main-OIMs-paper], each oscillator can be parameterized by a phase $\phi_{i}\in[0,2\pi]$ whose dynamics follow $\frac{d\phi_{i}}{dt^{\prime}}=-\sum_{j=1,\ j\neq i}^{n}J_{ij}\sin(\phi_{i}-\phi_{j})-h_{i}\sin(\phi_{i})-S_{i}\sin(2\phi_{i})$.
@@
These dynamics are equivalent to gradient descent $\frac{d\phi_{i}}{dt^{\prime}}=-\frac{\partial V}{\partial\phi_{i}}$, for an energy function [^1] Symmetric $J_{ij}$ includes $\frac{1}{2}$ to avoid double counting. [basharNoteAnalyzingStability2023]: $V=-\frac{1}{2}\sum_{i,j\ i\neq j}^{n}J_{ij}\cos(\phi_{i}-\phi_{j})-\sum_{i}^{n}h_{i}\cos(\phi_{i})-\sum_{i}^{n}\frac{S_{i}}{2}\cos(2\phi_{i}).$
@@ 3 Equilibrium Propagation
Equilibrium Propagation (EP) [scellierEquilibriumPropagationBridging2017, laborieuxScalingEquilibriumPropagation2020] is a local alternative to backpropagation for networks whose dynamics are characterized by the convergence to a stationary point, avoiding a separate backward pass by nudging these steady states toward desired outputs.
@@
The updates to the trainable parameters $\{\theta\}$ are then given by [laborieuxScalingEquilibriumPropagation2020] $\Delta\theta=\eta\hat{\nabla}^{\rm EP}(\beta)$ for learning rate $\eta$ and where:
@@
This update effectively steers the free stationary point toward configurations with lower loss by decreasing energy at $s^{\beta}_{*}$ and increasing it at $s^{-\beta}_{*}$. In the limit of infinitesimal $\beta$, this update equals gradient descent on the loss [laborieuxScalingEquilibriumPropagation2020]:
@@
Moreover, during the nudged phase dynamics (before convergence), the instantaneous EP updates match Backpropagation Through Time (BPTT) in the $\beta\to 0$ limit [werbosBackpropagationTimeWhat1990]:
@@ 4 Implementation on Oscillator Ising Machines
We implement a dense MLP neural network on an OIM with $n_{x}$ input neurons (fixed during dynamics), $n_{h}$ hidden neurons, and $n_{y}$ output neurons. Each non-input neuron corresponds to an oscillator, with the dynamical state $\phi=(\phi^{(h)},\phi^{(y)})$, and trainable parameters ${\theta}=({w_{ij}^{(x,h)}},{w_{ij}^{(h,y)}},{b_{i}^{(h)}},{b_{i}^{(y)}})$.
@@
Training follows standard EP. We use He weight initialization[heDelvingDeepRectifiers2015] with zero biases. For each input $x$: (i) Configure the OIM parameters as specified above; (ii) Initialize all oscillators to the reference state [^2] This corresponds to setting all neuron activations $\cos(\frac{\pi}{2})=0$. We found that any random initialization of phases also works well, provided the same random reference state is used consistently across all dynamics. $\phi_{0}=\{\frac{\pi}{2}\}$; (iii) Run the free phase ($\beta=0$) dynamics until convergence to $\phi_{*}$; (iv) Run the positive and negative nudged phase dynamics (initialized at $\phi_{*}$) until convergence to $\phi_{*}^{+\beta}$ and $\phi_{*}^{-\beta}$.
@@ 5 Results
We validate our OIM–EP implementation on MNIST [lecunGradientbasedLearningApplied1998], Fashion-MNIST [xiaoFashionMNISTNovelImage2017], and the MNIST/100 subset (1,000/100 train/test used in prior Ising machine benchmarks [laydevantTrainingIsingMachine2024]).
Training used PyTorch [^3] Code available at https://github.com/alexgower/OIM-Equilibrium-Propagation. with explicit Euler integration of the OIM dynamics. For full MNIST/FMNIST we used $T=4000$ free-phase steps and $K=400$ nudged steps with time step $\epsilon=0.45$; for MNIST/100 we used $T=3500$, $K=350$, $\epsilon=0.5$. We used constant nudging factors $\beta=0.1$ (full MNIST/FMNIST) and $\beta=0.05$ (MNIST/100) unless otherwise stated. The parameters $(T,K,\epsilon,\beta)$ were selected by increasing the free-phase duration $\epsilon T$ and decreasing $\beta$ until the symmetric EP updates closely matched BPTT updates (EP–BPTT inset of Fig. 1) while maintaining stable dynamics, and by choosing the nudged-phase duration $\epsilon K$ large enough that the nudged updates saturated. We used fixed-length trajectories and verified that $\max_{i}|d\phi_{i}/dt^{\prime}|$ was negligible at the end of each free phase, indicating convergence. We used a batch size of 128 for full MNIST/FMNIST and 20 for MNIST/100, and layer-specific learning rates $(\eta^{(h)}_{w},\eta^{(y)}_{w},\eta^{(h)}_{b},\eta^{(y)}_{b})=(0.01,0.001,0.001,0.001).$
@@
OIM–EP MLPs with a 784–120–10 architecture achieve 90.6±1.7% on MNIST/100 (92.0±0.3% with noise), and a 784–500–10 MLP reaches 97.2±0.1%/88.0±0.1% on full MNIST/FMNIST (Fig. 1). These results surpass D-Wave annealers [laydevantTrainingIsingMachine2024] and match BPTT on the same OIMs (96.8±0.1% on MNIST), indicating accuracy is architecture-limited rather than training-limited. On FMNIST, OIMs outperform p-bit Ising machines [niaziTrainingDeepBoltzmann2024] and are competitive with other neuromorphic MLPs [zhangTemporalSpikeSequence2020].
@@
Overall, OIMs maintain accuracy under 10-bit parameters, 4-bit phase detection [graberIntegratedCoupledOscillator2024], and moderate phase noise, confirming viability for CMOS implementations [royExperimentsOscillatorBased2025].
@@ 6 Discussion
Physical OIMs promise dramatic speedups. With GHz-frequency oscillators reaching microsecond timescales for parameter updates and phase measurements [main-OIMs-paper], full MNIST training (50 epochs, 60k examples, 3 EP phases) could complete in seconds to minutes, compared to $\sim$40 hours for EP and $\sim$60 hours for BPTT in our simulations—an acceleration of several orders of magnitude. This speed and efficiency could make energy-based learning practical for tasks previously limited by training time.
@@
Acknowledgements
This work was supported by UKRI/EPSRC CASE Award 220191 in partnership with Nokia UK Limited, and by a G-Research academic grant, which helped support the presentation of this work at NeurIPS 2025.